# An Adaptive Gradient Regularization Method

### v3

@@ Abstract
Optimizer plays an important role in neural network training with high efficiency and performance. Weight update based on its gradient is the central part of the optimizer. It has been shown that normalization and standardization operation on weight and gradient can accelerate the training process and improve performance such as Weight Standardization (WS), weight normalization (WN) and gradient normalization (GN); there is also gradient centralization (GC). In this work, we introduce a new optimization technique based on the gradient magnitude in a gradient vector named adaptive gradient regularization (AGR), which normalizes the gradient vector in all dimensions as a coefficient vector and subtracts the product of the gradient and its coefficient vector by the vanilla gradient. It can be viewed as an adaptive gradient clipping method. We show that the AGR can improve the loss function Lipschitzness with a more stable training process and better generalization performance. AGR is very simple to be embedded into vanilla optimizers such as Adan and AdamW with only three lines of code. Our experiments are conducted in image generation, image classification and language representation, which shows that our AGR improves the training result. We have released our code in https://github.com/hjiang24/AGR-method/tree/master
@@ 1.    Introduction
Deep learning has achieved great success in broad fields, e.g., vision tasks (Guo et al., 2022), natural language processing (Goldberg, 2016), and recommendation system (Zhang et al., 2019b ). Apart from the increasing development of large datasets (Yu et al., 2015) and abundant computing resources, e.g., GPUs and TPUs, elegant neural network architectures (Liu et al., 2017) and optimization algorithms (Ghadimi and Lan, 2013) are also the key ingredients in deep learning. For optimization algorithms, stochastic gradient-based algorithms contribute to the success of deep learning on large-scale datasets with satisfactory speed and local minimum loss.
@@
There are a variety of stochastic gradient-based algorithms, among which SGD (Ruder, 2016; Bottou, 2012) is the earliest and most representative algorithm for its simplicity and efficiency in many sophisticated network architectures. Subsequently, many variants of SGD have been proposed for faster training speed and better prediction accuracy, e.g. SGD with momentum (SGDM) (Qian, 1999), Adam (Kingma and Ba, 2014), AdamW (Loshchilov and Hutter, 2017), Adagrad (Duchi et al., 2011), RMSProp (Zou et al., 2019) and ACProp (Zhuang et al., 2021). The introduction of first-order momentum and second-order momentum combined with adaptive learning strategy make them popular in current deep neural networks (DNNs) training.
There are some techniques to improve training stability. The most prevalent method is normalization, e.g., batch normalization (BN) (Ioffe and Szegedy, 2015), group normalization (GN) (Wu and He, 2018), layer normalization (LN) (Ba et al., 2016), and instance normalization (IN) (Ulyanov et al., 2016). The design on sample/feature, weight, and weight gradient are useful and general optimization techniques on DNN training, which is to find a satisfactory solution in weight space with speed. BN utilizes statistics of samples in a mini-batch to approximate statistics of entire training sets to solve the internal covariate shift problem. Hence, training speed has been improved significantly under the same accuracy with much higher learning rates and less care about initialization.
@@
Many Network architectures have incorporated BN as the indispensable layer, and the ideology it invents provides insights for subsequent research. Weight Normalization (WN) (Salimans and Kingma, 2016) inspired by BN decouples the length of weight vectors from their direction, which is simple and provides much of the speed-up of full batch normalization. Weight Standardization (WS) (Qiao et al., 2019) applies zero mean and unit variance on the weights to smooth the loss landscape by reducing the Lipschitz constants of the loss and the gradients. Inspired by WN and WS, gradient centralization (GC) (Yong et al., 2020a ) is proposed to operate on the gradients of weight by simply centralizing the gradient vectors to have zero mean, which could regularize both the weight space and output feature space to boost the generalization performance of DNNs.
Inspired by the research mentioned above, we propose a new optimization technique to operate on weight gradients by Hadamard product with learnable coefficients instead of the zero-mean standardization of weight gradients, which is simple and could also be a kind of adaptive gradient regularization (AGR). The main contributions of our work are as follows:
  • We propose a new optimizer technique, named Adaptive Gradient Regularization (AGR), which could not only smooth the training loss but also have better generalization performance in multiple network architectures on different tasks, e.g., image generation, image classification, and token prediction.
  • We analyze the theoretical foundation of AGR and find it operates a new regularization method on gradients and provides a direction in the weight iteration process, which is also viewed as a new adaptive learning rate scaling method.
@@ 2.    Related work
Weight: Weight Normalization (WN) (Salimans and Kingma, 2016; Gitman and Ginsburg, 2017) is a simple reparameterization of the weight vectors in DNNs that accelerates the convergence of SGD optimization, which is an alternative to Batch Normalization (BN). Weight Standardization (WS) (Qiao et al., 2019) is targeted at the micro-batch training where the small batch sizes are not enough for training DNNs with Batch Normalization (BN), which normalizes weight vectors instead of features like Batch Normalization and can smooth the loss landscape by reducing the Lipschitz constants of the loss and the gradients.
@@
Gradient: Gradient is calculated for loss back-propagation in deep network training to update weight. Various gradient improvement methods have been proposed for a more stable training process and better generalization performance. Momentum is defined as the first-order momentum and the second-order momentum of gradients to be introduced as optimizer development method (Qian, 1999; Botev et al., 2017). In SGDM, momentum helps SGD accelerate in the relevant direction and dampen oscillations by considering the previous updates (Ruder, 2016). SGD with Nesterov Acceleration (Nesterov, 2014; Tran et al., 2022) takes into account the momentum term in the calculation of the gradient, by considering the future position of the parameter. Gradient clipping (Zhang et al., 2019a ; Chen et al., 2020; Menon et al., 2019) and gradient normalization (Chen; Wu et al., 2021) were proposed for training convergence acceleration. Gradient clipping is to solve the gradient exploding problem by scaling the gradient to keep it small when the gradient becomes too large, improving training stability and accelerating convergence. gradient normalization is similar to weight normalization decoupling the length of gradients from their direction. Gradient Centralization (GC) (Yong et al., 2020b ) centralizes gradients to have zero means, which can be viewed as a projected gradient descent method with a constrained loss function and regularizes the weight space and output feature space boosting the generalized performance of DNNs.
@@
Distinguished from weights and activations, the distribution of gradients is approximately lognormal(Chmiel et al., 2020; Guo et al., 2021). The distribution has a few gradients with huge magnitude and many gradients with small magnitude. The logarithmic scale of such distribution is close to a symmetric normal (Gaussian) distribution around the mean as zero, which is consistent with the assumption that gradient in the neural network is sampled from Gaussian distribution (Wiedemann et al., 2020). During the training process, the gradient distribution should be maintained as unchanged as possible. In other words, the difference between gradients needs to be as small as possible. That is why gradient clipping was proposed to accelerate DNN training. However, the current gradient modification methods could not adjust each gradient according to its magnitude dynamically. It is notis challenging to set an appropriate threshold in gradient clipping.
@@
Adaptive learning rate: Determining a good learning rate can prevent the system from diverging in terms of the loss function and slow learning (Chen et al., 2018a ). Adagrad (Duchi et al., 2011) scales the learning rate adaptively in each iteration for all dimensions based on the sum of the outer product of the gradients, considering all previous updates. However, the learning rate will be almost unchanged if the gradient is large at the beginning and small subsequently. RMSprop (Zou et al., 2019) introduces decay factors in the outer product of the gradients instead of simple summing of them, solving the problem of Adagrad. Adam (Kingma and Ba, 2014) and AdamW (Loshchilov and Hutter, 2017) also adapt the ideology to scale the learning rate adaptively. Adan (Xie et al., 2022) is an adaptive Nesterov algorithm reformulating the vanilla Nesterov momentum to develop a new Nesterov momentum estimation (NME), which improves the model training speed across multiple networks.
@@ 3    Adaptive Gradient Regularization
3.1 Motivation
Normalization and standardization have been applied to features and weight vectors, respectively, such as BN (Ioffe and Szegedy, 2015), WN (Salimans and Kingma, 2016; Gitman and Ginsburg, 2017), WS (Qiao et al., 2019), etc. These methods can reduce the Lipschitz constant of the loss function and smooth the optimization landscape. Apart from applying to samples and weight vectors, GC (Yong et al., 2020a ) computes the mean of gradients and centralizes the gradients to have zero mean instead of operating standardization on gradients, which also have good Lipschitz property, smooth the DNN training and improves the generalization performance with a strong restraint by projecting the gradient on a hyperplane. Apart from the statistics operation on gradients, how about applying adaptive factors to all gradients in each iteration? Based on an assumption we propose that when gradient magnitudes in all directions remain consistent, the training process will be more smooth and stable. Hence, we compute the factor by the ratio of each gradient to the sum of gradients in a neural network layer. Then, each gradient subtracts the product of itself and the corresponding factor to replace rigid traditional gradient clipping as a new regularization method. It can also be viewed as an adaptive learning rate scaling method.
3.2 Definitions and Preliminaries
There are some notations to be defined. We consider the gradients of the weight matrix in convolutional layers and fully connected layers. For convolutional layers, we denote weight matrix (kernel) $\textbf{W}_{\mathcal{C}}\in\mathbb{R}^{\mathcal{C}_{in}\times\mathcal{C}_{out}\times\mathcal{K}\times\mathcal{K}}$, $\mathcal{C}_{in}$ is the number of input channels, $\mathcal{C}_{out}$ is the number of output channels, $\mathcal{K}\times\mathcal{K}$ is the size of convolutional kernel. For fully connected layers, we denote weight matrix $\textbf{W}_{\mathcal{F}}\in\mathbb{R}^{\mathcal{C}_{in}\times\mathcal{C}_{out}}$, $\mathcal{C}_{in}$ is the number of input channels, $\mathcal{C}_{out}$ is the number of output channels. For simplicity, we denote the weight matrix $\textbf{W}\in\mathbb{R}^{\mathcal{M}\times\mathcal{N}}$ in unified format. In addition, $\textbf{w}_{i,j}$ is the element in i-th row and j-th column in $\textbf{W}(i=1,2,...M;j=1,2,...N)$. Denote W the weight matrix and $\mathcal{L}$ the objective function. $\nabla_{\textbf{w}}\mathcal{L}$ and $\nabla_{\textbf{w}_{i,j}}\mathcal{L}$ denote the gradient of $\mathcal{L}$ w.r.t. the weight matrix W and weight $\textbf{w}_{i,j}$. $\left|\nabla_{\textbf{w}}\mathcal{L}\right|$ and $\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|$ denote the absolute value of $\nabla_{\textbf{w}}\mathcal{L}$ and $\nabla_{\textbf{w}_{i,j}}\mathcal{L}$. $\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}$ is the sum of $\nabla_{\textbf{w}_{i,j}}\mathcal{L}$ in all dimensions.
@@ 3.3 AGR Formulation
For the fully connected layer and convolutional layer, suppose the gradient of weight $\textbf{w}_{i,j}$ is $\nabla_{\textbf{w}_{i,j}}\mathcal{L}(i=1,2,...M;j=1,2,...N)$. Denote the AGR operator $\Psi$, defined as below:
$\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})=\nabla_{\textbf{w}_{i,j}}\mathcal{L}-\alpha_{i,j}\nabla_{\textbf{w}_{i,j}}\mathcal{L}$ (1)
$\alpha_{i,j}=\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}$ (2)
The simple calculation $\alpha_{i,j}$is the ratio of the gradient absolute value to the sum of the gradient absolute values of the gradient matrix in all dimensions. Hence, the weight update in each iteration with AGR is below:
$\textbf{w}^{t+1}=\textbf{w}^{t}-\eta\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})$ (3)
3.4 Apply AGR to AdamW/Adan Optimizers
It is convenient and compact to embed ARG into AdamW (Loshchilov and Hutter, 2017) and Adan (Xie et al., 2022) with just two lines of code. In Algorithm 1 and Algorithm 2, we update the weight by replacing $\nabla_{\textbf{w}_{i,j}}\mathcal{L}$ with $\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})$ to implement ARG in AdamW and Adan, respectively. We also find no additional computational cost and training time across networks. Notably, we only apply the AGR to the gradients and the first-order momentum instead of the second-order momentum because we expect to reduce the learning rate and gradient consistently when the gradient is large.

### v4

# Adaptive Gradient Regularization: A Faster and Generalizable Optimization Technique for Deep Neural Networks

@@ Abstract
Stochastic optimization plays a crucial role in the advancement of deep learning technologies. Over the decades, significant effort has been dedicated to improving the training efficiency and robustness of deep neural networks, via various strategies including gradient normalization (GN) and gradient centralization (GC). Nevertheless, to the best of our knowledge, no one has considered to capture the optimal gradient descent trajectory, by adaptively controlling gradient descent direction. To address this concern, this paper is the first attempt to study a new optimization technique for deep neural networks, using the sum normalization of a gradient vector as coefficients, to dynamically regularize gradients and thus to effectively control optimization direction. The proposed technique is hence named as the adaptive gradient regularization (AGR). It can be viewed as an adaptive gradient clipping method. The theoretical analysis reveals that the AGR can effectively smooth the loss landscape, and hence can significantly improve the training efficiency and model generalization performance. We note that AGR can greatly improve the training efficiency of vanilla optimizers’ including Adan and AdamW, by adding only three lines of code. The final experiments conducted on image generation, image classification, and language representation, demonstrate that the AGR method can not only improve the training efficiency but also enhance the model generalization performance.
@@ 1.    Introduction
Over the decades, deep learning has achieved great success in promoting technological and scientific advancements in broad fields including computer vision (He et al. 2016a; LeCun, Bengio, and Hinton 2015), natural language processing (Devlin et al. 2018; Vaswani 2017), scientific computing (Wang et al. 2023), and etc. On the one hand, the wide development of large-scale datasets (Deng et al. 2009) and abundant computing resources including GPUs and TPUs, greatly benefit the advancement of deep learning technologies. On the other hand, elegant neural network architectures (Liu et al. 2017) and stochastic gradient descent algorithms (Bottou 2012), ensure deep neural networks to be able to capture large-scale datasets, with satisfying training efficiency and excelllent learning performance.
@@
Recent studies on the optimization techniques for neural networks, aim to accelerate the neural network training efficiency and also to improve the model generalization performance. To the end, there has been a growing number of stochastic gradient-based algorithms, among which stochastic gradient descent (SGD) (Ruder 2016; Bottou 2012) is the earliest and most representative one for its simplicity and efficiency in training many sophisticated network architectures. Afterward, many variants of SGD have been dedicated to improving training efficiency and learning performance, e.g. SGD with momentum (SGDM) (Qian 1999), Adam (Kingma and Ba 2014), AdamW (Loshchilov and Hutter 2017), Adagrad (Duchi, Hazan, and Singer 2011), RMSProp (Tieleman 2012) and ACProp (Zhuang et al. 2021). In particular, the introduction of first-order and second-order momentum combined with adaptive learning strategies, enables us to effectively train deep neural networks (DNNs).
@@
Notably, some recent studies (Chen et al. 2018b; Qiao et al. 2019) find that the statistics of gradients can be utilized to regularize gradients, and hence to stabilize the training of neural networks. The most prevalent method is normalization, e.g., batch normalization (BN) (Ioffe and Szegedy 2015), group normalization (GN) (Wu and He 2018), layer normalization (LN) (Ba, Kiros, and Hinton 2016), and instance normalization (IN) (Ulyanov, Vedaldi, and Lempitsky 2016). BN utilizes the statistics of samples in a mini-batch to approximate the statistics of the entire training set, to circumvent the internal covariate shift issue. As a result, the training speed can be significantly improved, without sacrificing the learning performance. BN has been incorporated into the design of many network architectures, and its ideology greatly benefits the subsequent studies (Salimans and Kingma 2016; Huang et al. 2017; Qiao et al. 2019). For instance, inspired by BN, weight normalization (WN) (Salimans and Kingma 2016) decouples the length of weight vectors from their direction, and hence can effectively speed up the full batch normalization in training deep neural networks. Weight standardization (WS) (Qiao et al. 2019) regularizes the weights to have zero mean and unit variance, to smooth the optimization landscape by reducing the Lipschitz constants of the loss function and gradients. Inspired by WN and WS, gradient centralization (GC) (Yong et al. 2020a) is proposed to operate on the gradients of weight, by simply centralizing the gradient vectors to have zero mean, which can regularize both the weight space and ambient feature space, to boost the generalization performance of DNNs.
Despite making some progress in accelerating neural network training efficiency, these optimization techniques still fail to adaptively control gradient descent direction, which may lead to finding the sub-optimal gradient descent trajectories. To mitigate this gap, we thoroughly study a novel optimization technique to adjust the gradient descent direction dynamically. To the end, we adaptively regularize the gradients using the sum normalization of a gradient vector as regularization coefficients, and hence can better approximate the optimal gradient descent trajectory. More specifically, we provide both the theoretical analysis and extensive experimental results to validate our hypothesis and the efficacy of the developed optimization technique. The main contributions of our work are as follows:
  • A novel gradient regularization strategy, is proposed to dynamically control the gradient descent direction.
  • A new optimization technique is developed by adaptively regularizing gradients on the element-scale, to improve the training efficiency and model generalization performance for deep neural networks. We hence call it the Adaptive Gradient Regularization (AGR). In particular, theoretical analysis are provided to guarantee that the AGR can effectively accelerate the training efficiency and improve the model generalization performance for deep neural networks, by smoothing the loss landscape.
  • Extensive experiments conducted on various tasks including image generation, image classification, and language representation, demonstrate the superior performance of the proposed AGR, in terms of improving the training efficiency for deep neural networks.
@@ 2.    Related Work
Weight: Weight Normalization (WN)  (Salimans and Kingma 2016; Gitman and Ginsburg 2017) is a simple reparameterization of the weight vectors in DNNs that accelerates the convergence of SGD optimization, which is an alternative to Batch Normalization (BN). Weight Standardization (WS) (Qiao et al. 2019) is targeted at the micro-batch training where the small batch sizes are not enough for training DNNs with Batch Normalization (BN), which normalizes weight vectors instead of features like Batch Normalization and can smooth the loss landscape by reducing the Lipschitz constants of the loss and the gradients.
@@
Gradient: Gradient is calculated for loss back-propagation in deep network training to update weight. Various gradient improvement methods have been proposed for a more stable training process and better generalization performance. Momentum is defined as the first-order momentum and the second-order momentum of gradients to be introduced as optimizer development method (Qian 1999; Botev, Lever, and Barber 2017). In SGDM, momentum helps SGD accelerate in the relevant direction and dampen oscillations by considering the previous updates (Ruder 2016). SGD with Nesterov Acceleration (Nesterov 2014; Tran, Scheinberg, and Nguyen 2022) takes into account the momentum term in the calculation of the gradient, by considering the future position of the parameter. Gradient clipping (Zhang et al. 2019; Chen, Wu, and Hong 2020; Menon et al. 2019) and gradient normalization (Chen et al. 2018b; Wu et al. 2021) were proposed for training convergence acceleration. Gradient clipping is to solve the gradient exploding problem by scaling the gradient to keep it small when the gradient becomes too large, improving training stability and accelerating convergence. gradient normalization is similar to weight normalization decoupling the length of gradients from their direction. Gradient Centralization (GC) (Yong et al. 2020b) centralizes gradients to have zero means, which can be viewed as a projected gradient descent method with a constrained loss function and regularizes the weight space and output feature space boosting the generalized performance of DNNs.
@@
Distinguished from weights and activations, the distribution of gradients is approximately lognormal(Chmiel et al. 2020; Guo, Liu, and Zhang 2021). The distribution has a few gradients with huge magnitude and many gradients with small magnitude. The logarithmic scale of such distribution is close to a symmetric normal (Gaussian) distribution around the mean as zero, which is consistent with the assumption that gradient in the neural network is sampled from Gaussian distribution (Wiedemann et al. 2020). During the training process, the gradient distribution should be maintained as unchanged as possible. In other words, the difference between gradients needs to be as small as possible. That is why gradient clipping was proposed to accelerate DNN training. However, the current gradient modification methods could not adjust each gradient according to its magnitude dynamically. It is challenging to set an appropriate threshold in gradient clipping.
@@
Adaptive learning rate: Determining a good learning rate can prevent the system from diverging in terms of the loss function and slow learning (Chen et al. 2018a). Adagrad (Duchi, Hazan, and Singer 2011) scales the learning rate adaptively in each iteration for all dimensions based on the sum of the outer product of the gradients, considering all previous updates. However, the learning rate will be almost unchanged if the gradient is large at the beginning and small subsequently. RMSprop (Tieleman 2012) introduces decay factors in the outer product of the gradients instead of simply summing them, solving the problem of Adagrad. Adam (Kingma and Ba 2014) and AdamW (Loshchilov and Hutter 2017) also adapt the ideology to scale the learning rate adaptively. Adan (Xie et al. 2022) is an adaptive Nesterov algorithm reformulating the vanilla Nesterov momentum to develop a new Nesterov momentum estimation (NME), which improves the model training speed across multiple networks.
In this paper, we propose a gradient regularization technique in the element scale based on their magnitude in an adaptive way.
@@ 3    Adaptive Gradient Regularization
Normalization and standardization have been applied on features and weight vectors, respectively, to speed up neural network training, such as BN (Ioffe and Szegedy 2015), WN (Salimans and Kingma 2016; Gitman and Ginsburg 2017), and WS (Qiao et al. 2019), etc. These methods can reduce the Lipschitz constant of the loss function and smooth the optimization landscape. Apart from applying to samples and weight vectors, GC (Yong et al. 2020a) computes the mean of gradients, and then centralizes the gradients to have zero mean, instead of performing standardization on gradients. GC can effectively improve the Lipschitzness of the loss function, and thus can improve the generalization performance, with a strong restraint by projecting the gradient on a hyperplane. The statistical operation on gradients by gradient normalization or centralization in multi-dimensions, does not consider the gradient magnitude on the element scale, and thus distributes the fixed coefficient to all the gradients. It may lead to capturing sub-optimal gradient descent trajectories since we cannot dynamically control the optimization direction. To address this limitation, we introduce the adaptive coefficient for each gradient by the sum normalization of the gradient vector in a neural network layer. Then, each gradient subtracts the product of itself and the corresponding coefficient, to adaptively approximate the optimal gradent descent direction.
3.1 Notations
We consider the gradients of the weight matrix in convolutional layers and fully connected layers. For convolutional layers, we denote the weight matrix (kernel) by $\textbf{W}_{\mathcal{C}}\in\mathbb{R}^{\mathcal{C}_{in}\times\mathcal{C}_{out}\times\mathcal{K}\times\mathcal{K}}$, in which $\mathcal{C}_{in}$ is the number of input channels, $\mathcal{C}_{out}$ is the number of output channels, and $\mathcal{K}\times\mathcal{K}$ is the size of convolutional kernel. For fully connected layers, we denote the weight matrix by $\textbf{W}_{\mathcal{F}}\in\mathbb{R}^{\mathcal{C}_{in}\times\mathcal{C}_{out}}$, in which $\mathcal{C}_{in}$ is the number of input channels, and $\mathcal{C}_{out}$ is the number of output channels. For simplicity, we denote the weight matrix $\textbf{W}\in\mathbb{R}^{\mathcal{M}\times\mathcal{N}}$ in an unified format. In addition, $\textbf{w}_{i,j}$ is the $(i,j)$-th entry of W. Denote W the weight matrix, and $\mathcal{L}$ the objective function. Hence, $\nabla_{\textbf{w}}\mathcal{L}$ and $\nabla_{\textbf{w}_{i,j}}\mathcal{L}$ denote the gradient of $\mathcal{L}$ w.r.t. the weight matrix W and the weight $\textbf{w}_{i,j}$, respectively. We use $\left|\nabla_{\textbf{w}}\mathcal{L}\right|$ to denote the absolute value of $\nabla_{\textbf{w}}\mathcal{L}$. Let $\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}$ to be the sum of $\nabla_{\textbf{w}_{i,j}}\mathcal{L}$ over all the dimensions.
@@ 3.2 AGR Formulation
For the fully connected layer and convolutional layer, suppose the gradient of weight $\textbf{w}_{i,j}$ is $\nabla_{\textbf{w}_{i,j}}\mathcal{L}(i=1,2,...M;j=1,2,...N)$. We denote the AGR operator by $\Psi$ and give the definition as below:
$\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})=\nabla_{\textbf{w}_{i,j}}\mathcal{L}-\alpha_{i,j}\nabla_{\textbf{w}_{i,j}}\mathcal{L},$ (1)
$\alpha_{i,j}=\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\sum_{i,j}\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}.$ (2)
The adaptive coefficient $\alpha_{i,j}$ is the ratio of the absolute value of the gradient to the sum of the absolute values of the gradients over all the dimensions.This method considers the gradient magnitude in the element scale, and thus can assign an adaptive coefficient to each gradient, to ensure adaptive gradient regularization. The weight update in each iteration using AGR is
$\textbf{w}^{t+1}=\textbf{w}^{t}-\eta\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L}),$ (3)
where $\eta$ is the learning rate.
3.3 Applying AGR to AdamW/Adan Optimizers
It is convenient to embed AGR into AdamW (Loshchilov and Hutter 2017) and Adan (Xie et al. 2022) with just three lines of code. More specifically, we update the weights by replacing $\nabla_{\textbf{w}_{i,j}}\mathcal{L}$ with $\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})$ to add AGR into AdamW and Adan as shown in Algorithm 1 and Algorithm 2, respectively. We also found no additional computational cost and training time in multiple neural networks. Notably, we only apply the AGR to the gradients and the first-order momentum instead of the second-order momentum because we expect to reduce the learning rate and regularize the gradient consistently when the gradient is large.